OpenStudy (vladthepro):

C. What is the solution to the equation1 over the square root of 8 = 4(m + 3)? m = − 15 over 4 m = − 11 over 4 m = 5 over 4 m = 9 over 4

OpenStudy (faiqraees):

square both sides

OpenStudy (kelsey.benton00):

\[\frac{ 1 }{ \sqrt{8} }=4(m+3)\] Does the equation look like that?

OpenStudy (vladthepro):

yes

OpenStudy (kelsey.benton00):

Ok, then you need to square both sides, as @FaiqRaees said, and then what do you get?

OpenStudy (vladthepro):

how would i do that?

OpenStudy (kelsey.benton00):

\[(\frac{ 1 }{ \sqrt{8} })^{2}=4(m+3)^{2}\]

OpenStudy (kelsey.benton00):

That, in turn, gives you \[\frac{ 1 }{ 8 }=4(m+3)^{2}\]

OpenStudy (kelsey.benton00):

What does \[(m+3)^{2}\] equal? Use the FOIL method.

OpenStudy (vladthepro):

2m+6 = 3/2

OpenStudy (kelsey.benton00):

How did you get that answer? Do you know what the FOIL method is?

OpenStudy (vladthepro):

Can that be explained, plz? ;-;

OpenStudy (anonymous):

\[m=\frac{1}{16} \left(\sqrt{2}-48\right) \]It appears that all of the answer options are incorrect.

OpenStudy (vladthepro):

Hmm.. interesting